OpenStudy (anonymous):

differentiate y= 2 sec x - csc x

OpenStudy (anonymous):

Here's what i've got so far.... \[y=2secxtanx - cscxcotx\]

OpenStudy (anonymous):

\[y'=[2\sec(x)\tan(x)]-[-\csc(x)\cot(x)] \rightarrow y'=2\sec(x)\tan(x)+\csc(x)\cot(x)\] @meaghan716

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You can leave it at that or simplify further by making tan(x), sin(x)/cos(x), and do the same for sec, csc, cot to whatever they're equal to. After all that you get:\[y'=\frac{ 2\sin^{3}(x)+\cos^3(x) }{ \cos^2(x) \sin^2(x) }\] @meaghan716

OpenStudy (anonymous):

Ok thank you sooo much